# IRAD: Implicit Representation-driven Image Resampling against Adversarial Attacks

### v1

@@ Abstract
We introduce a novel approach to counter adversarial attacks, namely, image resampling. Image resampling transforms a discrete image into a new one, simulating the process of scene recapturing or rerendering as specified by a geometrical transformation. The underlying rationale behind our idea is that image resampling can alleviate the influence of adversarial perturbations while preserving essential semantic information, thereby conferring an inherent advantage in defending against adversarial attacks. To validate this concept, we present a comprehensive study on leveraging image resampling to defend against adversarial attacks. We have developed basic resampling methods that employ interpolation strategies and coordinate shifting magnitudes. Our analysis reveals that these basic methods can partially mitigate adversarial attacks. However, they come with apparent limitations: the accuracy of clean images noticeably decreases, while the improvement in accuracy on adversarial examples is not substantial. We propose implicit representation-driven image resampling (IRAD) to overcome these limitations. First, we construct an implicit continuous representation that enables us to represent any input image within a continuous coordinate space. Second, we introduce SampleNet, which automatically generates pixel-wise shifts for resampling in response to different inputs. Furthermore, we can extend our approach to the state-of-the-art diffusion-based method, accelerating it with fewer time steps while preserving its defense capability. Extensive experiments demonstrate that our method significantly enhances the adversarial robustness of diverse deep models against various attacks while maintaining high accuracy on clean images.
@@ 1 Introduction
Adversarial attacks can mislead powerful deep neural networks by adding optimized adversarial perturbations to clean images (Croce & Hein, 2020b; Kurakin et al., 2018; Goodfellow et al., 2014; Guo et al., 2020), posing severe threats to intelligent systems. Existing works enhance the adversarial robustness of deep models by retraining them with the adversarial examples generated on the fly (Tramèr et al., 2018; Shafahi et al., 2019; Andriushchenko & Flammarion, 2020) or removing the perturbations before processing them (Liao et al., 2018; Huang et al., 2021b; Ho & Vasconcelos, 2022; Nie et al., 2022). These methods assume that the captured image is fixed. Nevertheless, in the real world, the observer can see the scene of interest several times via different observation ways since the real world is a continuous space and allows observers to resample the signal reflected from the scene. This could benefit the robustness of the perception system. We provide an illustrative example in Fig. 1 (a): when an image taken from a particular perspective is subjected to an attack that misleads the deep model (e.g., WRN28-10), by altering the viewing way, the same object can be re-captured and correctly classified. Such a process is also known as image resampling (Dodgson, 1992) that transforms a discrete image into a new one, simulating the process of scene recapturing or rerendering as specified by a geometrical transformation. In this work, we aim to study using image resampling to enhance the adversarial robustness of deep models against adversarial attacks.
@@ 2 Related Work
Adversarial Attack and Defense. White box attacks assume the attacker has full knowledge of the target model, including its architecture, weights, and hyperparameters. This allows the attacker to generate adversarial examples with high fidelity using gradient-based optimization techniques, such as FGSM (Goodfellow et al., 2014), BIM (Kurakin et al., 2018), PGD (Madry et al., 2017), Square Attack (Andriushchenko et al., 2020), and transferability-based attacks (Liu et al., 2016). AutoAttack (Croce & Hein, 2020b) has been proposed as a more comprehensive evaluation framework for adversarial attacks. AutoAttack combines several white box and black box attacks into a single framework and evaluates the robustness of a model against these attacks.
Adversarial defense can be categorized into two main types: adversarial training and adversarial purification (Nie et al., 2022). Adversarial training involves incorporating adversarial samples during the training process (Goodfellow et al., 2014; Madry et al., 2017; Athalye et al., 2018; Rade & Moosavi-Dezfooli, 2021; Ding et al., 2018), and training with additional data generated by generative models (Sehwag et al., 2021). On the other hand, adversarial purification functions as a separate defense module during inference and does not require additional training time for the classifier (Guo et al., 2017; Xu et al., 2017; Sun et al., 2019; Ho & Vasconcelos, 2022).
@@ 3.2 Naive Implementation
where $\mathcal{U}(0,\gamma)$ is a uniform distribution with the minimum and maximum being 00 and $\gamma$, respectively (See 4th column in Fig. 2). We can set different ranges $\gamma$ to see the changes in adversarial robustness. We can use the two reconstruction methods with different sampling strategies against adversarial attacks. We take the CIFAR10 dataset and the WideResNet28-10 (Zagoruyko & Komodakis, 2016) as examples. We train the WideResNet28-10 on CIFAR10 dataset and calculate the clean testing dataset’s accuracy, also known as the standard accuracy (SA).
@@
Then, we conduct the AutoAttack (Croce & Hein, 2020b) against the WideResNet28-10 on all testing examples and calculate the accuracy, denoted as the robust accuracy (RA). We employ image resampling to process the input, which can be the clean image or adversarial example, and the processed input is fed to the WideResNet28-10. Then, we can calculate the SA and RA of WideResNet28-10 with image resampling. We evaluate the effectiveness of the two naive reconstruction methods, i.e., bilinear interpolation and nearest interpolation with the sampling strategies defined in Eq. (7) and Eq. (8), which are denoted as $\textsc{IR}(\text{bil},\textsc{Sampler}(d\leavevmode\nobreak\ \text{or}\leavevmode\nobreak\ \gamma))$ and $\textsc{IR}(\text{nea},\textsc{Sampler}(d\leavevmode\nobreak\ \text{or}\leavevmode\nobreak\ \gamma)$, respectively. More details about the dataset, the model architecture, and the adversarial attack AutoAttack are deferred to Sec. 6.
@@ 3.3 Discussions and Motivations
Based on the findings presented in Table 1, the following observations emerge:
@@
❶ Bilinear interpolation with SI-sampling (i.e., $\textsc{IR}(\text{bil},\textsc{Sampler}(d=1.5))$) significantly enhances robust accuracy, albeit at the expense of a modest reduction in standard accuracy. In contrast, using nearest interpolation with SI-sampling results in only marginal variations in SA and RA. ❷ For both interpolation methods, SV-sampling leads to notable increases in RA, while simultaneously causing a significant reduction in SA. Overall, we see some effectiveness of leveraging naive image resampling methods against adversarial attacks. Nevertheless, such methods are far from being able to achieve high standard and robust accuracy at the same time. The reasons are that the naive interpolation-based reconstruction methods could not remove the perturbations while the sampling strategies are not designed to preserve the semantic information. As the example shown in Fig. 2, we feed the adversarial example to the bilinear interploration-based method with SI-sampling and SV-sampling, respectively, and use the Grad-CAM to present the semantic variations before and after resampling. Clearly, the two sampling strategy do not preserve the original semantic information properly. To address the issues, we propose a novel IR method in Sec. 4.
@@ 4.2 Implicit Representation
  • Super-resolution. We can also train the model via the super-resolution task as done in Chen et al. (2021). That is, given a low-resolution input $\mathbf{I}_{\text{lr}}$ that is downsampled from a clean image $\mathbf{I}$, we aim to generate a higher resolution by sampling more coordinates and feeding them to Eq. (9), thus we can get a large-size image $\hat{\mathbf{I}}_{\text{hr}}$. We can also use the loss function $L_{1}(\mathbf{I},\hat{\mathbf{I}}_{\text{hr}})$ to train the model.
@@ 4.4 Implementation Details
We follow the recent work (Chen et al., 2021) and set the deep model in Lim et al. (2017) as the $\textsc{Feat}(\cdot)$ to extract the embedding of the $\mathbf{I}$. We set $\varphi(\cdot)$ and the SampleNet as a five-layer MLP, respectively. We utilize the adversarial denoising task (See discussion in Sec. 6.3 ) to train the models through a two-stage training strategy. Specifically, we first train the implicit representation (i.e., $\textsc{Feat}(\cdot)$ and $\varphi(\cdot)$) on the training dataset, and then we fix their weights and train the SampleNet for sampling. Note that we follow a black-box setup; that is, we train our model based on adversarial examples crafted from ResNet18 and test the effectiveness on other deep models (See experimental section). Please refer to the supplementary material for other training details.
@@ 5 Relationship and Extension to SOTAs
Relationship to implicit representation-based method (e.g., (Ho & Vasconcelos, 2022)). Ho & Vasconcelos (2022) employ implicit representation (Chen et al., 2021) to remove the adversarial perturbation and can enhance the robust accuracy under attacks significantly while preserving the high accuracy on the clean images. Different to Ho & Vasconcelos (2022), we employ the implicit representation (Chen et al., 2021) as a part of image resampling and study the influences of different training tasks. More importantly, our method contains the SampleNet that can automatically predict the suitable pixel-wise shifting and further recover the semantic information. We demonstrate the effectiveness of our method over (Ho & Vasconcelos, 2022) in Sec. 6.1.
@@
Table 2: Comparison on CIFAR10, CIFAR100 and ImageNet via AutoAttack ($\mathcal{\epsilon_{\infty}}=8/255$). "–" indicates the absence of a corresponding pre-trained model in the original paper.
@@
w.o. Defense 94.77 0 47.39 81.04 3.46 42.25 76.72 0 38.36
Bit Reduction (Xu et al., 2017) 92.66 1.05 46.86 77.40 3.74 40.57 73.74 1.88 37.81
Jpeg (Dziugaite et al., 2016) 83.66 50.79 67.23 63.67 40.86 52.27 73.28 33.96 53.62
Randomization (Xie et al., 2017) 93.87 6.86 50.37 76.22 9.76 42.99 74.04 19.81 46.93
Median Filter 79.66 42.54 61.10 68.56 35.79 52.18 71.66 17.59 44.63
@@
IRAD 91.70 89.72 90.71 75.32 71.76 73.54 72.14 71.60 71.87
Table 3: Comparison on CIFAR10, CIFAR100 and ImageNet via BPDA. "–" indicates the absence of a corresponding pre-trained model in the original paper.
@@
w.o. Defense 94.77 0.02 47.40 80.45 0.40 40.43 76.72 0.00 38.36
Bit Reduction (Xu et al., 2017) 92.66 0.40 46.53 77.28 1.29 39.29 73.74 0.00 36.87
Jpeg (Dziugaite et al., 2016) 83.65 5.50 44.58 63.62 12.32 37.97 73.28 0.04 36.66
Randomization (Xie et al., 2017) 94.05 34.81 64.43 76.10 33.02 54.56 74.06 27.92 50.99
Median Filter 79.66 25.12 52.39 68.07 29.24 48.66 71.66 0.02 35.84
STL (Sun et al., 2019) 90.65 2.10 46.38 72.62 0.00 37.14
DISCO (Ho & Vasconcelos, 2022) 89.25 22.60 55.93 73.76 28.20 50.98 72.44 0.34 36.39
DiffPure (Nie et al., 2022) 89.15 87.06 88.11 68.85 61.42 65.13
IRAD 91.70 74.32 83.01 74.04 42.71 58.37 72.14 71.12 71.63
@@ 6 Experimental Results
Datasets. During training and evaluation of IRAD, We use three datasets: CIFAR10 (Krizhevsky et al., a), CIFAR100 (Krizhevsky et al., b) and ImageNet (Deng et al., 2009). To make training datasets of the same size, we randomly select 50000 images in ImageNet, which includes 50 samples for each class. IRAD is trained on pairs of adversarial and clean training images generated by the PGD attack (Madry et al., 2017) on ResNet18 (He et al., 2016). The PGD attack uses an $\epsilon$ value of 8/255 and 100 steps, with a step size 2/255. The SampleNet is also trained using adversarial-clean pairs generated by the PGD attack. In this part, we mainly present part of the results; other results will be shown in the supplementary materials.
Attack scenarios.Oblivious adversary scenario: We follow setups in RobustBench (Croce et al., 2021) and use the AutoAttack as the main attack method for the defense evaluation since AutoAttack is an ensemble of several white-box and black-box attacks, including two kinds of PGD attack, the FAB attack (Croce & Hein, 2020a), and the square attack (Andriushchenko et al., 2020), allowing for a more comprehensive evaluation.
@@
In addition to AutoAttack, we also report the results against FGSM (Goodfellow et al., 2014), BIM (Kurakin et al., 2018), PGD (Madry et al., 2017), RFGSM (Tramèr et al., 2018), TPgd (Zhang et al., 2019b), APgd (Croce & Hein, 2020b), EotPgd (Liu et al., 2018), FFgsm (Wang et al., 2020), MiFgsm (Dong et al., 2018), and Jitter (Schwinn et al., 2021). We evaluate methods by utilizing adversarial examples generated through victim models that are pre-trained on clean images. ❷ Adaptive adversary scenario: We consider a challenging scenario where attackers know both the defense methods and classification models. Compared to the oblivious adversary, this is particularly challenging, especially regarding test-time defense (Athalye et al., 2018; Sun et al., 2019; Tramer et al., 2020). To evaluate the performance, we utilize the BPDA method (Athalye et al., 2018) that can circumvent defenses and achieve a high attack success rate, particularly against test-time defenses. ❸ AutoAttack-based Adaptive adversary scenario. We regard the IRAD and the target model as a whole and use AutoAttack to attack the whole process.
Baelines. We compare with 7 representative testing-time adversarial defense methods as shown in Table 2 including 2 SOTA methods, i.e., DISCO (Ho & Vasconcelos, 2022) and DiffPure (Nie et al., 2022). Note that, we run all methods by ourselves with their released models for a fair comparison. We also report more comparisons with training-time methods in the supplementary material.

### v3

@@ Abstract
We introduce a novel approach to counter adversarial attacks, namely, image resampling. Image resampling transforms a discrete image into a new one, simulating the process of scene recapturing or rerendering as specified by a geometrical transformation. The underlying rationale behind our idea is that image resampling can alleviate the influence of adversarial perturbations while preserving essential semantic information, thereby conferring an inherent advantage in defending against adversarial attacks. To validate this concept, we present a comprehensive study on leveraging image resampling to defend against adversarial attacks. We have developed basic resampling methods that employ interpolation strategies and coordinate shifting magnitudes. Our analysis reveals that these basic methods can partially mitigate adversarial attacks. However, they come with apparent limitations: the accuracy of clean images noticeably decreases, while the improvement in accuracy on adversarial examples is not substantial. We propose implicit representation-driven image resampling (IRAD) to overcome these limitations. First, we construct an implicit continuous representation that enables us to represent any input image within a continuous coordinate space. Second, we introduce SampleNet, which automatically generates pixel-wise shifts for resampling in response to different inputs. Furthermore, we can extend our approach to the state-of-the-art diffusion-based method, accelerating it with fewer time steps while preserving its defense capability. Extensive experiments demonstrate that our method significantly enhances the adversarial robustness of diverse deep models against various attacks while maintaining high accuracy on clean images. We released our code in https://github.com/tsingqguo/irad.
@@ 1 Introduction
Adversarial attacks can mislead powerful deep neural networks by adding optimized adversarial perturbations to clean images (Croce & Hein, 2020b; Kurakin et al., 2018; Goodfellow et al., 2014; Guo et al., 2020; Huang et al., 2023a), posing severe threats to intelligent systems. Existing works enhance the adversarial robustness of deep models by retraining them with the adversarial examples generated on the fly (Tramèr et al., 2018; Shafahi et al., 2019; Andriushchenko & Flammarion, 2020) or removing the perturbations before processing them (Liao et al., 2018; Huang et al., 2021b; Ho & Vasconcelos, 2022; Nie et al., 2022). These methods assume that the captured image is fixed. Nevertheless, in the real world, the observer can see the scene of interest several times via different observation ways since the real world is a continuous space and allows observers to resample the signal reflected from the scene. This could benefit the robustness of the perception system. We provide an illustrative example in Fig. 1 (a): when an image taken from a particular perspective is subjected to an attack that misleads the deep model (e.g., ResNet50), by altering the viewing way, the same object can be re-captured and correctly classified. Such a process is also known as image resampling (Dodgson, 1992) that transforms a discrete image into a new one, simulating the process of scene recapturing or rerendering as specified by a geometrical transformation. In this work, we aim to study using image resampling to enhance the adversarial robustness of deep models against adversarial attacks.
@@ 2 Related Work
Adversarial Attack and Defense. White box attacks assume the attacker has full knowledge of the target model, including its architecture, weights, and hyper-parameters. This allows the attacker to generate adversarial examples with high fidelity using gradient-based optimization techniques, such as FGSM (Goodfellow et al., 2014), BIM (Kurakin et al., 2018), PGD (Madry et al., 2017), and others (Huang et al., 2023b). Other attacks also include black box attacks like Square Attack (Andriushchenko et al., 2020) and patch-wise attacks (Gao et al., 2020), as well as transferability-based attacks (Liu et al., 2016; Wang & He, 2021; Wang et al., 2021). AutoAttack (Croce & Hein, 2020b) has been proposed as a more comprehensive evaluation framework for adversarial attacks. AutoAttack combines several white box and black box attacks into a single framework and evaluates the robustness of a model against these attacks.
Adversarial defense can be categorized into two main types: adversarial training and adversarial purification (Nie et al., 2022). Adversarial training involves incorporating adversarial samples during the training process (Goodfellow et al., 2014; Madry et al., 2017; Athalye et al., 2018; Rade & Moosavi-Dezfooli, 2021; Ding et al., 2018; Zhang et al., 2020a; Jia et al., 2022b), and training with additional data generated by generative models (Sehwag et al., 2021). On the other hand, adversarial purification functions as a separate defense module during inference and does not require additional training time for the classifier (Guo et al., 2017; Xu et al., 2017; Sun et al., 2019; Ho & Vasconcelos, 2022).
@@ 3.2 Naive Implementation
where $\mathcal{U}(0,\gamma)$ is a uniform distribution with the minimum and maximum being 00 and $\gamma$, respectively (See 4th column in Fig. 2). We can set different ranges $\gamma$ to see the changes in adversarial robustness.
@@
We can use the two reconstruction methods with different sampling strategies against adversarial attacks. We take the CIFAR10 dataset and the WideResNet28-10 (Zagoruyko & Komodakis, 2016) as examples. We train the WideResNet28-10 on CIFAR10 dataset and calculate the clean testing dataset’s accuracy, also known as the standard accuracy (SA). Then, we conduct the AutoAttack (Croce & Hein, 2020b) against the WideResNet28-10 on all testing examples and calculate the accuracy, denoted as the robust accuracy (RA). We employ image resampling to process the input, which can be the clean image or adversarial example, and the processed input is fed to the WideResNet28-10. Then, we can calculate the SA and RA of WideResNet28-10 with image resampling. We evaluate the effectiveness of the two naive reconstruction methods, i.e., bilinear interpolation and nearest interpolation with the sampling strategies defined in Eq. (7) and Eq. (8), which are denoted as $\textsc{IR}(\text{bil},\textsc{Sampler}(d\leavevmode\nobreak\ \text{or}\leavevmode\nobreak\ \gamma))$ and $\textsc{IR}(\text{nea},\textsc{Sampler}(d\leavevmode\nobreak\ \text{or}\leavevmode\nobreak\ \gamma)$, respectively. More details about the dataset, the model architecture, and the adversarial attack AutoAttack are deferred to Sec. 6.
@@ 3.3 Discussions and Motivations
Based on the findings presented in Table 1, the following observations emerge: ❶ Bilinear interpolation with SI-sampling (i.e., $\textsc{IR}(\text{bil},\textsc{Sampler}(d=1.5))$) significantly enhances robust accuracy, albeit at the expense of a modest reduction in standard accuracy. In contrast, using nearest interpolation with SI-sampling results in only marginal variations in SA and RA. ❷ For both interpolation methods, SV-sampling leads to notable increases in RA, while simultaneously causing a significant reduction in SA. Overall, we see some effectiveness of leveraging naive image resampling methods against adversarial attacks. Nevertheless, such methods are far from being able to achieve high standard and robust accuracy at the same time. The reasons are that the naive interpolation-based reconstruction methods could not remove the perturbations while the sampling strategies are not designed to preserve the semantic information. As the example shown in Fig. 2, we feed the adversarial example to the bilinear interploration-based method with SI-sampling and SV-sampling, respectively, and use the Grad-CAM to present the semantic variations before and after resampling. Clearly, the two sampling strategy do not preserve the original semantic information properly. To address the issues, we propose a novel IR method in Sec. 4.
@@ 4.2 Implicit Representation
  • Super-resolution. We can also train the model via the super-resolution task as done in (Chen et al., 2021). That is, given a low-resolution input $\mathbf{I}_{\text{lr}}$ that is downsampled from a clean image $\mathbf{I}$, we aim to generate a higher resolution by sampling more coordinates and feeding them to Eq. (9), thus we can get a large-size image $\hat{\mathbf{I}}_{\text{hr}}$. We can also use the loss function $L_{1}(\mathbf{I},\hat{\mathbf{I}}_{\text{hr}})$ to train the model.
@@ 4.4 Implementation Details
We follow the recent work (Chen et al., 2021) and set the deep model in (Lim et al., 2017) as the $\textsc{Feat}(\cdot)$ to extract the embedding of the $\mathbf{I}$. We set $\varphi(\cdot)$ and the SampleNet as a five-layer MLP, respectively. We utilize the adversarial denoising task (See discussion in Sec. 6.3 ) to train the models through a two-stage training strategy. Specifically, we first train the implicit representation (i.e., $\textsc{Feat}(\cdot)$ and $\varphi(\cdot)$) on the training dataset, and then we fix their weights and train the SampleNet for sampling. Note that we follow a black-box setup; that is, we train our model based on adversarial examples crafted from ResNet18 and test the effectiveness on other deep models (See experimental section). Please refer to the Appendix for other training details.
@@ 5 Relationship and Extension to SOTAs
Relationship to implicit representation-based method (e.g., (Ho & Vasconcelos, 2022)). (Ho & Vasconcelos, 2022) employ implicit representation (Chen et al., 2021) to remove the adversarial perturbation and can enhance the robust accuracy under attacks significantly while preserving the high accuracy on the clean images. Different to (Ho & Vasconcelos, 2022), we employ the implicit representation (Chen et al., 2021) as a part of image resampling and study the influences of different training tasks. More importantly, our method contains the SampleNet that can automatically predict the suitable pixel-wise shifting and further recover the semantic information. We demonstrate the effectiveness of our method over (Ho & Vasconcelos, 2022) in Sec. 6.1.
@@
Table 2: Comparison on CIFAR10, CIFAR100 and ImageNet via AutoAttack ($\mathcal{\epsilon_{\infty}}=8/255$ for CIFAR10 and CIFAR100, $\mathcal{\epsilon_{\infty}}=4/255$ for ImageNet). "–" indicates no corresponding pre-trained model in the original paper.
@@
w.o. Defense 94.77 0 47.39 81.66 3.48 42.57 76.72 0 38.36
Bit Reduction (Xu et al., 2017) 92.66 1.05 46.86 74.47 6.56 40.52 73.74 1.88 37.81
Jpeg (Dziugaite et al., 2016) 83.66 50.79 67.23 60.87 38.36 49.62 73.28 33.96 53.62
Randomization (Xie et al., 2017) 93.87 6.86 50.37 78.7 10.25 44.48 74.04 19.81 46.93
Median Filter 79.66 42.54 61.10 57.32 31.18 44.25 71.66 17.59 44.63
NRP (Naseer et al., 2020) 92.89 3.82 48.36 77.17 12.67 44.92 72.52 20.40 46.46
@@
IRAD 91.70 89.72 90.71 76.01 72.49 74.25 72.14 71.60 71.87
Table 3: Comparison on CIFAR10, CIFAR100, and ImageNet via BPDA.
@@
w.o. Defense 94.77 0.02 47.40 81.67 0.49 41.08 76.72 0.00 38.36
Bit Reduction (Xu et al., 2017) 92.66 0.40 46.53 74.47 0.62 37.55 73.74 0.00 36.87
Jpeg (Dziugaite et al., 2016) 83.65 5.50 44.58 60.87 5.79 33.33 73.28 0.04 36.66
Randomization (Xie et al., 2017) 94.05 34.81 64.43 78.85 21.03 49.94 74.06 27.92 50.99
Median Filter 79.66 25.12 52.39 57.32 11.52 34.42 71.66 0.02 35.84
NRP (Naseer et al., 2020) 92.89 0.27 52.39 77.17 0.46 38.82 72.52 0.02 36.27
STL (Sun et al., 2019) 90.65 2.10 46.38 72.62 0.02 36.32
DISCO (Ho & Vasconcelos, 2022) 89.25 22.60 55.93 72.58 16.90 44.74 72.44 0.34 36.39
DiffPure (Nie et al., 2022) 89.15 87.06 88.11 68.85 61.42 65.14
IRAD 91.70 74.32 83.01 76.00 62.78 69.39 72.14 71.12 71.63
@@ 6 Experimental Results
Datasets. During training and evaluation of IRAD, We use three datasets: CIFAR10 (Krizhevsky et al., a), CIFAR100 (Krizhevsky et al., b) and ImageNet (Deng et al., 2009). To make training datasets of the same size, we randomly select 50000 images in ImageNet, which includes 50 samples for each class. IRAD is trained on pairs of adversarial and clean training images generated by the PGD attack (Madry et al., 2017) on ResNet18 (He et al., 2016). The PGD attack uses an $\epsilon$ value of 8/255 and 100 steps, with a step size 2/255. The SampleNet is also trained using adversarial-clean pairs generated by the PGD attack. In this part, we mainly present part of the results; other results will be shown in the Appendix.
Attack scenarios.Oblivious adversary scenario: We follow setups in RobustBench (Croce et al., 2021) and use the AutoAttack as the main attack method for the defense evaluation since AutoAttack is an ensemble of several white-box and black-box attacks, including two kinds of PGD attack, the FAB attack (Croce & Hein, 2020a), and the square attack (Andriushchenko et al., 2020), allowing for a more comprehensive evaluation. In addition to AutoAttack, we also report the results against FGSM (Goodfellow et al., 2014), BIM (Kurakin et al., 2018), PGD (Madry et al., 2017), RFGSM (Tramèr et al., 2018), TPgd (Zhang et al., 2019b), APgd (Croce & Hein, 2020b), EotPgd (Liu et al., 2018), FFgsm (Wang et al., 2020), MiFgsm (Dong et al., 2018), and Jitter (Schwinn et al., 2021). We evaluate methods by utilizing adversarial examples generated through victim models that are pre-trained on clean images. ❷ Adaptive adversary scenario: We consider a challenging scenario where attackers know both the defense methods and classification models. Compared to the oblivious adversary, this is particularly challenging, especially regarding test-time defense (Athalye et al., 2018; Sun et al., 2019; Tramer et al., 2020). To evaluate the performance, we utilize the BPDA method (Athalye et al., 2018) that can circumvent defenses and achieve a high attack success rate, particularly against test-time defenses. ❸ AutoAttack-based Adaptive adversary scenario. We regard the IRAD and the target model as a whole and use AutoAttack to attack the whole process.
Baelines. We compare with 7 representative testing-time adversarial defense methods as shown in Table 2 including 2 SOTA methods, i.e., DISCO (Ho & Vasconcelos, 2022) and DiffPure (Nie et al., 2022). Note that, we run all methods by ourselves with their released models for a fair comparison. We also report more comparisons with training-time methods in the Appendix.